# Beyond Prompt Content: Enhancing LLM Performance via Content-Format Integrated Prompt Optimization

### v2

@@ 4.1 Experimental Setups
Dataset and Models. To rigorously evaluate CFPO, we selected a diverse set of tasks and models. Our benchmark tasks span various domains and complexities, including:
@@
  • Reasoning: GSM8K (Cobbe et al., 2021) and MATH500 (Hendrycks et al., 2021; Lightman et al., 2023) which require complex mathematical reasoning abilities.
  • Multiple-choice: ARC-Challenge (Clark et al., 2018), demanding understanding and selection among alternatives.
  • Classification: The Implicatures task from the Big-Bench benchmark (bench authors, 2023) to evaluate classification proficiency.
Our model selection includes a mix of foundational and instruction-tuned models to understand the generalizability of our approach:
  • Foundational Models: Mistral-7B-v0.1 (Jiang et al., 2023) and LLaMA-3.1-8B (Meta, 2024b) represent pre-trained models.
  • Instruction-Tuned Models: LLaMA-3-8B-Instruct (Meta, 2024a) and Phi-3-Mini-Instruct (Microsoft, 2024) represent models specifically fine-tuned for instruction following.
Furthermore, we use GPT-4 (2024-05-01-preview) as the LLM optimizer for content mutation and format generation (OpenAI, 2024a).
Implementation Details. The training process involved $20$ iterative rounds, each consisting of content and format optimization. During content optimization, case-diagnosis and Monte Carlo sampling each generate 4 prompts per round. A set of 40 test cases is used, with 5 correct and incorrect cases leveraged for case-diagnosis. The number of prompt-structured components decreases progressively from 4 to 1, narrowing the search space over time to enhance efficiency. For format optimization, 4 UCT-selected formats and 4 newly generated formats are used to generate new prompts. The coefficient in the UCT selection process $\alpha$ is set to $31e-$ 3. Beam search, with a budget of 8, is employed during mutations to ensure effective exploration. Eval data sizes are configured as 50, 300, 500, and 500 for BigBench-Classification, MATH500, GSM8K, and ARC-Challenge, respectively. The best performing prompt on the evaluation set for each method was selected and reported on the test set.
@@
Baselines. To evaluate the effectiveness of CFPO, we compared againt several commonly used and popular baselines. GrIPS (Prasad et al., 2023) performs syntactic phrase-level edits in instruction, representing a non-LLM-based optimization approach. APE (Zhou et al., 2023) and ProTeGi (Pryzant et al., 2023) both employ LLM to optimize prompt content, but differ in mutation strategy. APE adopts an instruction induction approach, while ProTeGi leverages test cases feedback with LLM to guide the mutation process. SAMMO (Schnabel and Neville, 2024) introduces a structured framework that incorporates a preliminary format mutation strategy, which relies on random selection from a predefined format pool. This choice of baselines enables a comprehensive assessment of CFPO’s capabilities against various types of optimization approaches. All methods were evaluated using consistent experimental configurations to ensure a fair comparison.
Initial Prompts. To establish a reasonable starting point, we employed a single in-context example without any further instruction as the initial prompt for each model and task, except for GrIPS which requires an initial instruction. Chain-of-Thought examples were employed for the reasoning tasks. We also report common baseline prompts, including 8-shot for GSM8K and 4-shot for MATH500. A comprehensive list of our initial prompts is in Appendix C.
@@ 4.2 Main Results
Table 1 summarizes the performance of CFPO in comparison with several state-of-the-art methods across four datasets. The results highlight the superior performance of CFPO, significantly outperforming the baseline prompt as well as competing methods. We observed that pre-trained models exhibit greater sensitivity to prompt formatting, leading to substantial improvements when optimized by CFPO. Notably, optimized prompts for pre-trained models tend to be longer and incorporate more in-context examples, suggesting that these characteristics better align with the optimization needs of pre-trained models (see Appendix D.1). In contrast, instruction-tuned models display relatively more robust results and smaller gains, likely due to their inherent adaptability and generalization.
For the reasoning tasks, GSM8K and MATH, prompt optimization is especially impactful due to the sensitivity of these tasks to prompt structure. CFPO, which integrates unified content and format optimization, delivers significant performance gains. Specifically, the improvement for GSM8K is more evident compared to the more challenging MATH task, where the inherent complexity limits the magnitude of improvement. Moreover, feedback-based methods like ProTeGi, SAMMO, and CFPO, consistently outperform the other baselines because they leverage iterative feedback for prompt refinement. In contrast, GRIPS, which is limited to phrase-level mutations, exhibits marginal improvements. These results underline the effectiveness of the integrated optimization strategy adopted by CFPO. The selected optimal prompts discovered by our approach can be found in Appendix D.
@@ 4.3 Ablation Study
Impact of the Format Optimizer. CFPO incorporates a unique format optimization process, leveraging LLM for format generation and a UCT-based strategy for format selection. To evaluate its effectiveness, we evaluated two variations of our method: (1) CFPOc, which optimizes content while keeping format fixed, and (2) CFPOc+Format, which first optimizes content, then performs a separate format optimization step. Table 2 shows that both CFPOc and CFPOc+Format underperform compared to the full CFPO approach, highlighting the importance of the integrated content and format optimization approach. The need for a joint optimization process which addresses the interdependence of content and format is essential for prompt optimization.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
@@
BBC ProTeGi 81.00 82.00
CFPOc 85.00 85.00
CFPOc+Format 88.00 89.00
CFPO 90.00 91.00
[…]
Effectiveness of Format Generation. We compared the full CFPO approach against a variant that uses format from initial format pool without using LLM for generation. As presented in Table 4, CFPO with format generation consistently outperforms the baseline relying solely on the initial pool. These results demonstrate the effectiveness of the proposed format exploration mechanism in enhancing both the quality and diversity of prompts.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
@@
BBC w/o Format Gen 88.00 87.00
with Format Gen 90.00 91.00
Table 3: Impact of format generation during prompt optimization.
Effectiveness of Format Selection. We further evaluated our UCT-based format selection process, compared it to a random selection from the format pool and a greedy selection without exploration (using $\alpha=0$ in Eq. (2)). As presented in Table3, CFPO consistently achieves the best performance across all experimental settings, demonstrating the efficacy of the UCT-based selection strategy.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
@@
Random 85.00 87.00
BBH UCT($\alpha=0$) 86.00 88.00
UCT(ours) 90.00 91.00
Table 4: Impact of different format selection strategies during optimization.
Effectiveness of the Content Optimizer. As presented in Table 2, we include ProTeGi (Pryzant et al., 2023), a baseline that optimizes only the content. In contrast, our CFPOc, which incorporates structured prompting and integrates correct cases for diagnosis, achieves significant performance improvements, which highlights the effectiveness of our content optimization strategy.
@@ 5 Conclusion
This paper introduces Content-Format Integrated Prompt Optimization (CFPO), a innovative methodology that concurrently optimizes both prompt content and format. CFPO incorporates the Prompt Renderer and the Query Format within a structured prompt template. By leveraging distinct optimization strategies, CFPO discovers high-performing prompts that outperform content-only methods, addressing a critical gap in existing research. Our results demonstrate the substantial significant influence of format on LLM performance, underscoring the necessity of a joint optimization approach. These findings emphasize the importance of integrating content and format considerations in prompt engineering. CFPO represents a significant advancement, empowering developers to design effective and robust prompts and unlocking the full potential of LLMs across diverse applications.
@@ A.1 Meta-Prompt Header Setup
A.2 Format Generation
@@
A.3 Content Optimization
A.3.1 Case-diagnosis and Revision
[META PROMPT HEADER]
[INCORRECT CASES]
Meanwhile, this prompt gets the following examples correct:
[CORRECT CASES]
Please encapsulate each suggestion using the following format:
<START>
<Prompt segment: [Segment name]>
[Suggestion goes here]
<END>
[META PROMPT HEADER]
The existing [COMPONENT NAME] segment contains:
[CURRENT CONTENT FOR THE COMPONENT]
[GENERATED DIAGNOSES]
[META PROMPT HEADER]
The existing EXAMPLES segment contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
Here are some suggestions for enhancing the EXAMPLES segment:
[GENERATED DIAGNOSES]
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
The [NUMBER OF GENERATED EXAMPLES] revised EXAMPLES are:
A.3.2 Monte-Carlo Sampling
[META PROMPT HEADER]
[CURRENT CONTENT FOR THE COMPONENT]
The varied [COMPONENT NAME] segment is as follows:
[META PROMPT HEADER]
The existing EXAMPLE set contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
Appendix B Appendix: Examples of Generated Format
Here we select several format generated by GPT4 in CFPO process.
B.1 Query Format

### v3

@@ 4.1 Experimental Setups
Models. We selected four open-source Large Language Models (LLMs) as our primary evaluation targets, including two foundational models, Mistral-7B-v0.1 (Jiang et al., 2023) and LLaMA-3.1-8B (Meta, 2024b), as well as two instruction-tuned models, LLaMA-3-8B-Instruct (Meta, 2024a) and Phi-3-Mini-Instruct (Microsoft, 2024). For content mutation and format generation during the optimization process, we employed GPT-4 (2024-05-01-preview) (OpenAI, 2024a).
Datasets. Our evaluation benchmark was designed to comprehensively assess model performance across a range of task types and difficulty levels, emphasizing diverse query formats. Specifically, we employed the following tasks:
  • Classification: The Implicatures task from the Big-Bench benchmark (bench authors, 2023).
  • Multiple-choice: ARC-Challenge (Clark et al., 2018) task.
@@
  • Reasoning: GSM8K (Cobbe et al., 2021) and MATH500 (Hendrycks et al., 2021; Lightman et al., 2023) requiring complex reasoning abilities.
@@
Baselines. We compared CFPO with several commonly used and popular baselines. GrIPS (Prasad et al., 2023) performs syntactic phrase-level edits in instruction, representing a non-LLM-based optimization approach. APE (Zhou et al., 2023) and ProTeGi (Pryzant et al., 2023) both employ LLM to optimize prompt content, but differ in mutation strategy. APE adopts an instruction induction approach, while ProTeGi leverages test cases feedback with LLM to guide the mutation process. SAMMO (Schnabel and Neville, 2024) introduces a structured framework that incorporates a preliminary format mutation strategy, which relies on random selection from a predefined format pool. All methods were evaluated using consistent experimental configurations to ensure a fair comparison. We also report common baseline prompts for reasoning tasks, including 8-shot for GSM8K and 4-shot for MATH500.
Implementation Details. To evaluate the generated prompts, we sample subsets from the training split for each benchmark (sizes: 50, 500, 500, and 300 examples respectively). Beam search (budget of 8) is used during prompt mutations. Each experiment was capped at 20 iterations. Early stopping was implemented, halting the process if performance plateaus. The number of prompt components the LLM could modify decreased linearly from 4 to 1 over the iterations (see Appendix B.1). We start with a single in-context example without any further instruction as the initial prompt for each model and task, except for GrIPS which requires an initial instruction. Detailed parameter settings and procedure are in the Appendix B.
@@ 4.2 Main Results
Table 1 summarizes the performance of CFPO compared with several state-of-the-art methods across four datasets. The results highlight the superior performance of CFPO, significantly outperforming the baseline prompt and competing methods.
The effectiveness of CFPO is particularly evident when compared to methods like GRIPS, which relies solely on phrase-level mutations, yielding only marginal improvements. This highlights the importance of iterative feedback for effective prompt refinement, a feature shared by ProTeGi, SAMMO, and CFPO. Furthermore, CFPO incorporates integrated format optimization that contributes significantly to the performance gains, particularly for previously challenging tasks.
The benefits of CFPO are particularly pronounced in reasoning tasks, such as GSM8K and MATH, known for their sensitivity to prompt structure. While improvements are observed on both datasets, the impact is more noticeable on GSM8K compared to the more complex MATH dataset, suggesting that task difficulty can moderate the attainable performance gains from prompt optimization.
Beyond performance metrics, we further evaluate the stability of CFPO and the computational cost, as detailed in Appendix C. Examples of optimal prompts are detailed in Appendix E.
@@ 4.3 Ablation Study
Task Method Llama3.1 Llama3-Ins
BBC ProTeGi 81.00 82.00
CFPOf 83.00 86.00
CFPOc 85.00 85.00
CFPOc+f 88.00 89.00
CFPO 90.00 91.00
@@
Table 2: Performance comparison of the full CFPO pipeline against ablated variants with format-only (CFPOf), content-only (CFPOc), and sequential content-then-format (CFPOc+f) optimization. CFPOc also compared to ProTeGi (Pryzant et al., 2023).
Impact of the CFPO Pipeline Components. To understand the contribution of each component in CFPO, we compared the full CFPO approach against the following ablated variants: (1) CFPOf (format-only): optimizes the prompt format while holding content constant, (2) CFPOc (content-only): optimizes the prompt content while holding format constant, and (3) CFPOc+f (sequential): performs iterative content optimization followed by a separate, single-step format optimization. We also include ProTeGi (Pryzant et al., 2023) as a baseline. While CFPOc shares the goal of content optimization with ProTeGi, CFPOc incorporates correct cases in diagnosis and component-wise mutation as key inovations. Table 2 presents the results of this analysis. CFPOc significantly outperforms ProTeGi, demonstrating the effectiveness of our content optimization strategy. Furthermore, the ablated variants all underperform compared to the complete CFPO pipeline. These results underscore the interdependence of content and format in prompt optimization, highlighting the importance of joint optimization for best performance.
Task Method Llama3.1 Llama3-Ins
Impact of Format Generation
BBC w/o Format Gen 88.00 87.00
with Format Gen 90.00 91.00
@@
Different Format Selection Strategies
Random 85.00 87.00
BBC UCT($\alpha=0$) 86.00 88.00
UCT(ours) 90.00 91.00
@@
Table 3: Ablation of format generation and comparison of format selection strategies.
Analysis of Format Generation. We compared the full CFPO approach against variant that uses format from initial format pool without using LLM for generation. As shown in Table 3, the version with format generation consistently outperforms the variant using only the initial pool. This highlights the benefit of our format exploration mechanism in expanding the prompt space. (See Appendix D.2 for the ablation study with different format generation models.)
Analysis of Format Selection Strategies. We further examined the effectiveness of our UCT-based format selection strategy. We compared it against two baselines: random selection from the format pool and a greedy selection approach (equivalent to setting $\alpha=0$ in Eq. (1), disabling exploration). Table 3 shows that CFPO with UCT-based selection consistently achieves the best performance across all settings, demonstrating the efficacy of balancing exploration and exploitation in format searching.
Optimizer Mistral Llama3.1 Llama3-Ins Phi-3
Qwen2.5-14B 50.49 58.76 80.14 88.48
GPT-4 (CFPO) 53.22 63.38 80.74 89.16
Table 4: Performance of CFPO using different optimizer models.
Exploring Different Optimizer Models. While GPT-4 was used as the primary optimizer model for CFPO, we also investigated the feasibility of using a more accessible model , Qwen2.5-14B-Instruct. Table 4 presents the performance of CFPO with Qwen2.5-14B-Instruct as the optimizer. The results are encouraging, demonstrating that CFPO can be effectively adapted to smaller-sized open-sourced models, albeit with some performance trade-offs, increasing its potential for broader applicability.
@@ 5 Conclusion
This paper introduces Content-Format Integrated Prompt Optimization (CFPO), an innovative methodology that concurrently optimizes both prompt content and format. By leveraging a structured prompt template, CFPO decouples these elements, enabling integrated optimization and addressing a significant gap in current research that largely overlooks the critical influence of format. Our results demonstrate the substantial significant influence of format on LLM performance, underscoring the necessity of a joint optimization approach. These findings emphasize the importance of integrating content and format considerations in prompt engineering. CFPO represents a significant advancement, empowering developers to design effective and robust prompts and unlocking the full potential of LLMs across diverse applications.
Limitations While the proposed method demonstrates promising results, there are several limitations worth noting. First, the effectiveness of the approach is task- and model-dependent. While the method generates promising prompts for specific tasks and models, it may not generalize as effectively to others—particularly tasks that are less sensitive to prompt structure or models that already possess strong reasoning capabilities, thereby limiting its broader applicability. Moreover, the iterative nature of the optimization process, with multiple mutation strategies, introduces computational complexity, which could hinder scalability in resource-constrained environments. Finally, we acknowledge the inherent difficulty in definitively separating content and format, especially in real-world scenarios where structured data can be embedded within content. Future work will focus on addressing these limitations, including exploring the application of CFPO to complex agentic tasks and improving the stability and scalability of the optimization process.
@@ A.1 Meta-Prompt Header Setup
A.2 Content Optimization
A.2.1 Case-diagnosis and Revision
[META PROMPT HEADER]
[INCORRECT CASES]
Meanwhile, this prompt gets the following examples correct:
[CORRECT CASES]
Please encapsulate each suggestion using the following format:
<START>
<Prompt segment: [Segment name]>
[Suggestion goes here]
<END>
[META PROMPT HEADER]
The existing [COMPONENT NAME] segment contains:
[CURRENT CONTENT FOR THE COMPONENT]
[GENERATED DIAGNOSES]
[META PROMPT HEADER]
The existing EXAMPLES segment contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
Here are some suggestions for enhancing the EXAMPLES segment:
[GENERATED DIAGNOSES]
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
The [NUMBER OF GENERATED EXAMPLES] revised EXAMPLES are:
A.2.2 Monte-Carlo Sampling
[META PROMPT HEADER]
[CURRENT CONTENT FOR THE COMPONENT]
The varied [COMPONENT NAME] segment is as follows:
[META PROMPT HEADER]
The existing EXAMPLE set contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
A.3 Format Generation
@@
Appendix B Appendix: Experimental Setup
B.1 Hyperparameter in Optimization Procedure
During content optimization, case-diagnosis and Monte Carlo sampling each generate 4 prompts per round. A set of 40 test cases is used, with 5 correct and incorrect cases leveraged for case-diagnosis. During optimization, the number of prompt components (e.g., Task Instruction, Task Detail, Output Format, Few-shot Examples) the LLM could modify simultaneously was dynamically adjusted across all the iterations. The number of components that could be modified decreases linearly from 4 to 1 over the iterations, promoting broad exploration of the prompt space initially and fine-grained refinement in later stages.
For format optimization, 4 UCT-selected formats and 4 newly generated formats are used to generate new prompts. The coefficient in the UCT selection process $\alpha$ is set to $31e-$ 3.
Each experiment was capped at a maximum of 20 optimization iterations. An early stopping criterion was implemented, halting the process if the performance did not improve for a specified number of consecutive iterations.
B.2 Model Generation Parameters
Key generation parameters for both the LLM optimizer (GPT-4) and the target evaluation models are detailed below.
  • LLM Optimizer (GPT-4): We utilized the following settings to generate and refine prompts: top_p=1.0, max_tokens=4096, seed=42, and temperature=1.0.
[…]
B.3 Evaluation Metric
Performance across all tasks was evaluated using the exact match metric. For direct answer tasks (BigBench-Classification and ARC-Challenge), the model’s direct prediction was assessed for an exact string match with the ground truth. For Chain-of-Thought reasoning tasks (GSM8K and MATH500), the final numerical answer was extracted from the generated explanation and compared to the correct solution for an exact match. Finally, the best-performing prompt identified on the evaluation set for each optimization method is reported on the corresponding test set.
B.4 Initial Prompt
B.4.1 Big-Bench Classification
Prompt Renderer: Directly Joint
Query Format: Input-Output
Examples:
Input: Speaker 1: You do this often?’ Speaker 2: Its my first time.’
Output: no
{{Query placeholder}}
B.4.2 ARC-Challenge
Prompt Renderer: Directly Joint
Query Format: MultiChoice_QA
Here are some examples:
Choices:
A: It reduces the amount of carbon dioxide production
B: It reduces the production of oxygen
C: It decreases the greenhouse effect
D: It decreases pollutants in the air
Answer: B
{{Query placeholder}}
B.4.3 GSM8K
Prompt Renderer: Directly Joint
Query Format: QA
{{Query placeholder}}
B.4.4 MATH500
Prompt Renderer: Directly Joint
Query Format: Question-Answer
Here are some examples:
[…]
{{Query placeholder}}
Appendix C Appendix: CFPO Performance, Stability, and Cost Analysis
C.1 CFPO Performance on MMLU
To further evaluate the Content-Format Integrated Prompt Optimization (CFPO) approach, we assessed its performance on the MMLU benchmark. Specifically, we focused on high school history categories (including high_school_european_history, high_school_us_history and high_school_world _history) to evaluate CFPO’s effectiveness on knowledge-based tasks.
Table 5 summarizes the MMLU performance, showcasing improvements achieved with CFPO, which reinforce the generalizability of CFPO, extending its demonstrated benefits from reasoning tasks (as presented in the main paper) to knowledge-based tasks. This expanded validation will be integrated into the revised manuscript.
Task Method Llama3.1 Llama3-Ins
MMLU Baseline 78.84 82.03
CFPO 81.74 83.77
Table 5: Ablation of format generation and comparison of format selection strategies.
C.2 Impact of In-context Examples and Prompt Length
Figure 5: Overview of in-context examples and text lengths for optimized prompts on various tasks and models.
C.3 Stability and Convergence Analysis
Model CFPO Reproduced CFPO
Mistral-7B-v0.1 53.22 51.87 ± 0.22
Llama-3.1-8B 63.38 62.70 ± 0.59
Llama-3-8B-Instruct 80.74 80.67 ± 0.17
Phi-3-Mini-Instruct 89.16 88.89 ± 0.31
Table 6: Stability Analysis: Comparison of reported CFPO scores with reproduced results (4 runs each).
Figure 6: Convergence behavior of format and content optimization (Mistral-7B-v0.1 on GSM8K). Values indicate performance scores, with improvements from the previous sub-round in parentheses.
A crucial aspect of our framework is its stability. To quantify this, we performed multiple independent runs of CFPO and analyzed the variance in performance. Table 6 presents a comparison of the originally reported CFPO scores with the mean and standard deviation of scores obtained from four independent runs. The consistently low standard deviations across all models demonstrate the robustness and stability of our CFPO framework.
Furthermore, we examined the convergence behavior of content and format optimization. Preliminary analysis using Mistral-7B-v0.1 on GSM8K (Figure 6) indicates that format optimization may converge faster than content optimization. The performance gains in each sub-round (content or format optimization) are shown in parentheses. This suggests that the optimization schedule could be refined.
This difference in convergence speed could stem from the distinct optimization strategies: during content optimization, a diverse set of correct and incorrect examples is resampled from the training set in each round. In contrast, format optimization relies on generating variations of existing formats, leading to reduced diversity as the UCT score converges. We hypothesize that allocating more resources to format optimization in early rounds, and subsequently shifting focus to content optimization, could improve overall efficiency. This is an area for further investigation in future work.
C.4 Cost Analysis
Operation Tokens (I/O) API Calls Est. Cost
Case-diagnose 2k/0.2k 32 \$ 0.83
Apply Feedback 0.5k/0.1k 110 \$ 0.93
Gen Variation 0.5k/0.2k 81 \$0.89
Gen Format 0.5k/0.1k 2 \$0.016
Gen Format Code 0.6k/0.2k 2 \$0.024
Total (per Round) 159.6k/36.6k 237 \$2.69
Table 7: Average token usage, API calls, and estimated cost per round of CFPO optimization. Tokens (I/O): typical input/output tokens per round, API Calls: average API calls per round, Est. Cost: estimated cost in \$.
Task Avg. Cost to Tar. Avg. Time to Tar.
BBC \$14.10 5.24 h
ARC-C \$16.81 6.25 h
GSM8K \$28.25 10.50 h
MATH500 \$19.85 7.38 h
Table 8: Average cost and time required for CFPO to reach the reported target state-of-the-art performance on each benchmark.
An important consideration is the computational cost associated with CFPO. Table 7 breaks down the average token usage, API calls, and estimated cost per round of optimization for each operation within CFPO. These costs are calculated using current API pricing for the utilized LLMs.
Furthermore, CFPO incorporates an early stopping mechanism that terminates the optimization process when performance improvement plateaus. This prevents unnecessary iterations and reduces overall cost. Table 8 details the average cost and time required to reach the reported performance on each benchmark task. As demonstrated in Table 8, the average computational cost per task remains manageable.
Appendix D Appendix: Format Generation Examples and Ablation Study
D.1 Examples of Generated Format
Here we select several format generated by GPT-4 in CFPO process.
D.1.1 Query Format
Highlight_Separator_Case
ANSWER > C
Cascading_Statements
Options:
-A True, True
-B False, False
-C True, False
-D False, True
Answer: C